OpenStudy (anonymous):

how to solve(solving systems by substitution) x=6y-11 3x-2y=-1

OpenStudy (anonymous):

Solve one of the equations for one of the variables in terms of the other one. Then substitute the equivalent expression in for the variable you solved for into the other equation. Solve for the other variable. Then plug that value back in to either equation to find the value of the first variable.

OpenStudy (anonymous):

so, i have to isolate x on the 2nd equation?

OpenStudy (gw2011):

Since x=6y-11, substitute this into the second equation and you get the following: 3(6y-11)-2y=-1 18y-33-2y=-1 16y=32 y=2 Now substitute y=2 into the first equation to solve for x: x=(6)(2)-11 x=12-11 x=1 To prove the correctness that x=1 and y=2 you can substitute these values into each of the two equations.